OpenStudy (anonymous):

if my answer is this: x-7/-4 why does my professor want me to write it like this: -1/4x+74!!!!!

OpenStudy (agent0smith):

That doesn't look right... post more info. Those two expressions aren't even remotely equivalent.

OpenStudy (anonymous):

it is: For each of the following functions, write an equation for the inverse function y = f ^−1(x) . f (x) = −4x + 7

OpenStudy (agent0smith):

First change it to y = −4x + 7 then rearrange for x=...

OpenStudy (anonymous):

y-7=-4x/-4 = y-7/-4=x = x-7/-4=y

OpenStudy (anonymous):

that doesn't equal -1/4x+7/4 ?

OpenStudy (agent0smith):

(y-7)/-4=x ^^^Make sure you use brackets. And it does, you originally wrote 74, not 74. \[\large \frac{ x-7 }{ -4 } = \frac{ x }{ -4 } - \frac{ 7 }{ -4 }\]

OpenStudy (agent0smith):

*not 7/4

OpenStudy (anonymous):

with the negative 4 into the seven fip the minus sign to a plus then you have positive 7/4?

OpenStudy (agent0smith):

Yep and x/-4 is the same as (-1/4)x

OpenStudy (anonymous):

she just wants it written like that because it looks better then?

OpenStudy (anonymous):

=)

OpenStudy (agent0smith):

That's the slope-intercept form. And looks neater.